OpenStudy (anonymous):

Is sin2(x) + cos2(x) = 1 the same as sin(α2) + cos(α2) = 1?

OpenStudy (unklerhaukus):

\[\sin^2(x)+\cos^2(x)=1\]

OpenStudy (anonymous):

so is that the same as sin(α^2) + cos(α^2) = 1?

OpenStudy (unklerhaukus):

i dont think you second equation is necessarily true

OpenStudy (zzr0ck3r):

no it is not the same.

OpenStudy (lgbasallote):

in the first part what you're squaring are sine and cosine in the second part, what you're squaring are the angles therefore, different things are squared, so different results